# SymTFT out of equilibrium: from time crystals to braided drives and Floquet codes

### 28 Dec 2023 v1

@@ C.2 $\pi$ and zero modes in (braided) Ising drives
$U_{\text{F}}^{(\bm{1},\langle e\rangle),\sigma}=e^{-i\sum_{j\neq 1}h_{j}X_{j}}e^{-iJ_{1}Z_{L}\tilde{Z}_{\tilde{L}}Y_{1}}e^{-i\sum_{j\neq 1}\kappa_{j}Z_{j-1}Z_{j}},$
@@
By it involving only terms from Eqs. (5) and (6), we can describe $U_{\text{F}}^{(\bm{1},\mathcal{A}),\sigma}$ using free fermions in each of the four $S=P,\tilde{Z}_{\tilde{L}}$ eigenspaces (with eigenvalues $p$, $\tilde{z}$ and $s=p\tilde{z}$ as above). Turning to $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$, the line operator $\smash{\overline{W}}^{(b)}$ does not cause difficulties, since $\smash{\overline{W}}^{(b)}=P^{r_{m}}\tilde{Z}_{\tilde{L}}^{r_{e}}$ with $r_{m},r_{e}\in\{0,1\}$. For each $p,\tilde{z}$, the unitary $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ thus acts via quadratic fermion gates; we denote the corresponding operator by $U_{\text{F},p,\tilde{z}}^{(b,\mathcal{A}),\sigma}=p^{r_{m}}\tilde{z}^{r_{e}}U_{\text{F},s}^{(\bm{1},\mathcal{A}),\sigma}$.
As before, $c_{1}$ is absent, hence $\gamma_{1}=c_{1}$ is an exact zero mode of $U_{\text{F},s}^{(\bm{1},\mathcal{A}),\sigma}$. We find its pair using $e^{-i\sum_{j\neq 1}h_{j}X_{j}}=e^{\sum_{ij}q_{ij}^{X}c_{i}c_{j}}$, $e^{-i\sum_{j\neq 1}\kappa_{j}Z_{j-1}Z_{j}}=e^{\sum_{ij}q_{ij}^{Z}c_{i}c_{j}}$, $e^{-iJ_{1}Z_{L}\tilde{Z}_{\tilde{L}}Y_{1}}\to e^{\sum_{ij}q_{ij}^{1,s}c_{i}c_{j}}$ with $(2L-1)\times(2L-1)$ antisymmetric matrices $q_{ij}^{Z},\ q_{ij}^{X},\ q_{ij}^{1,s}$. Using the correspondence between an operator $\hat{Q}=e^{\sum_{jk}q_{jk}c_{j}c_{k}}$ and matrix $Q=e^{4q}$ in $\hat{Q}c_{n}\hat{Q}^{-1}=\sum_{m}Q_{mn}c_{m}$ Kitaev (2001), the sought pair is $\gamma_{2}^{(s)}=\sum_{j=2}^{2L}v_{j}^{(s)}c_{j}$, with $v^{(s)}$ now satisfying $e^{4q^{X}}e^{4q^{1,s}}e^{4q^{Z}}v^{(s)}=v^{(s)}$.
@@
By $q_{ij}^{Z},\ q_{ij}^{X},\ q_{ij}^{1,s}$ being odd dimensional antisymmetric, such a $v^{(s)}$ always exist and, sufficiently away from criticality (i.e., from $\bar{J}=\bar{h}$ for the typical values $\bar{J}$ of $J_{i}$ and $\bar{h}$ of $h_{j}$ Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Berdanier et al. (2018)) is unique. The value of $\bar{J}/\bar{h}$, however, sets the spatial profile of $v^{(s)}$, making it increasingly delocalized upon approaching $\bar{J}=\bar{h}$. We illustrate this via numerical simulations, shown in Fig. C1.
From $\gamma_{2}^{(s)}$ one gets the many-body operator $\gamma_{2}$ as in Eq. (7). For $U_{\text{F}}^{(\bm{1},\mathcal{A}),\sigma}$, the operators $\tilde{\gamma}_{1}=i\gamma_{1}\Gamma_{1}$ and $\tilde{\gamma}_{2}=i\gamma_{2}\Gamma_{1}$ are again exact zero modes. For $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ (with $\mathcal{A}=\langle e\rangle$, $b=m$ or vice versa), by $\smash{\overline{W}}^{(b)}\tilde{\gamma}_{j}=-\tilde{\gamma}_{j}\smash{\overline{W}}^{(b)}$ ($j=1,2$), these are both exact $\pi$ modes. Finally, for the braided drive $U_{\text{F},R}^{(b,\mathcal{A}),\sigma}=i\tilde{\gamma}_{1}U_{\text{F}}^{(b,\mathcal{A}),\sigma}$, the operator $\tilde{\gamma}_{1}$ is an exact $\pi$ mode, while, by $\tilde{\gamma}_{1}\tilde{\gamma}_{2}=-\tilde{\gamma}_{2}\tilde{\gamma}_{1}$, the operator $\tilde{\gamma}_{1}$ is an exact zero mode. Thus we have an unpaired $\pi$ and zero mode each, both exact.

### 28 Feb 2024 v2

@@ C.2 $\pi$ and zero modes in (braided) Ising drives
$U_{\text{F}}^{(\bm{1},\langle e\rangle),\sigma}=e^{-i\sum_{j\neq 1}h_{j}X_{j}}e^{-iJ_{1}Z_{L}\tilde{Z}_{L}Y_{1}}e^{-i\sum_{j\neq 1}\kappa_{j}Z_{j-1}Z_{j}},$
@@
By it involving only terms from Eqs. (5) and (6), we can describe $U_{\text{F}}^{(\bm{1},\mathcal{A}),\sigma}$ using free fermions in each of the four $S=P,\tilde{Z}_{L}$ eigenspaces (with eigenvalues $p$, $\tilde{z}$ and $s=p\tilde{z}$ as above). Turning to $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$, the line operator $\smash{\overline{W}}^{(b)}$ does not cause difficulties, since $\smash{\overline{W}}^{(b)}=P^{r_{m}}\tilde{Z}_{L}^{r_{e}}$ with $r_{m},r_{e}\in\{0,1\}$. For each $p,\tilde{z}$, the unitary $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ thus acts via quadratic fermion gates; we denote the corresponding operator by $U_{\text{F},p,\tilde{z}}^{(b,\mathcal{A}),\sigma}=p^{r_{m}}\tilde{z}^{r_{e}}U_{\text{F},s}^{(\bm{1},\mathcal{A}),\sigma}$.
As before, $c_{1}$ is absent, hence $\gamma_{1}=c_{1}$ is an exact zero mode of $U_{\text{F},s}^{(\bm{1},\mathcal{A}),\sigma}$. We find its pair using $e^{-i\sum_{j\neq 1}h_{j}X_{j}}=e^{\sum_{ij}q_{ij}^{X}c_{i}c_{j}}$, $e^{-i\sum_{j\neq 1}\kappa_{j}Z_{j-1}Z_{j}}=e^{\sum_{ij}q_{ij}^{Z}c_{i}c_{j}}$, $e^{-iJ_{1}Z_{L}\tilde{Z}_{L}Y_{1}}\to e^{\sum_{ij}q_{ij}^{1,s}c_{i}c_{j}}$ with $(2L-1)\times(2L-1)$ antisymmetric matrices $q_{ij}^{Z},\ q_{ij}^{X},\ q_{ij}^{1,s}$. Using the correspondence between an operator $\hat{Q}=e^{\sum_{jk}q_{jk}c_{j}c_{k}}$ and matrix $Q=e^{4q}$ in $\hat{Q}c_{n}\hat{Q}^{-1}=\sum_{m}Q_{mn}c_{m}$ Kitaev (2001), the sought pair is $\gamma_{2}^{(s)}=\sum_{j=2}^{2L}v_{j}^{(s)}c_{j}$, with $v^{(s)}$ now satisfying $e^{4q^{X}}e^{4q^{1,s}}e^{4q^{Z}}v^{(s)}=v^{(s)}$.
@@
By $q_{ij}^{Z},\ q_{ij}^{X},\ q_{ij}^{1,s}$ being odd dimensional antisymmetric, such a $v^{(s)}$ always exists and, sufficiently away from criticality (i.e., from $\bar{J}=\bar{h}$ for the typical values $\bar{J}$ of $J_{i}$ and $\bar{h}$ of $h_{j}$ Khemani et al. (2016); Else et al. (2016); von Keyserlingk and Sondhi (2016a, b); von Keyserlingk et al. (2016); Khemani et al. (2017, ); Berdanier et al. (2018)) is unique. The value of $\bar{J}/\bar{h}$, however, sets the spatial profile of $v^{(s)}$, making it increasingly delocalized upon approaching $\bar{J}=\bar{h}$. We illustrate this via numerical simulations, shown in Fig. C1.
From $\gamma_{2}^{(s)}$ one gets the many-body operator $\gamma_{2}$ as in Eq. (7). For $U_{\text{F}}^{(\bm{1},\mathcal{A}),\sigma}$, the operators $\tilde{\gamma}_{1}=i\gamma_{1}\Gamma_{1}$ and $\tilde{\gamma}_{2}=i\gamma_{2}\Gamma_{1}$ are again exact zero modes. For $U_{\text{F}}^{(b,\mathcal{A}),\sigma}$ (with $\mathcal{A}=\langle e\rangle$, $b=m$ or vice versa), by $\smash{\overline{W}}^{(b)}\tilde{\gamma}_{j}=-\tilde{\gamma}_{j}\smash{\overline{W}}^{(b)}$ ($j=1,2$), these are both exact $\pi$ modes. Finally, for the braided drive $U_{\text{F},R}^{(b,\mathcal{A}),\sigma}=i\tilde{\gamma}_{1}U_{\text{F}}^{(b,\mathcal{A}),\sigma}$, the operator $\tilde{\gamma}_{1}$ is an exact $\pi$ mode, while, by $\tilde{\gamma}_{1}\tilde{\gamma}_{2}=-\tilde{\gamma}_{2}\tilde{\gamma}_{1}$, the operator $\tilde{\gamma}_{2}$ is an exact zero mode. Thus we have an unpaired $\pi$ and zero mode each, both exact.